Find the area bounded by the curves y^2=4x and y=2x-4.


1
Expert's answer
2021-04-15T07:20:36-0400

Curve "y^2 = 4x" is parabola and equation "y=2x-4" represent the straight line.

These two curves will intersect at "(4,4) , (1,-2)" .

To find the intersection point, simply solve the equation, "(2x-4)^2 = 4x"

"4x^2+16-16x=4x"

"\implies x^2-5x+4=0"

"(x-1)(x-4)=0"

"x = 1, 4"

so "y=-2, 4"


Area under the curve is given by,

"A = \int_{-2}^{4} |(\frac{y+4}{2}) - \frac{y^2}{4} |dy"

"A = \int_{-2}^{0} |(\frac{y+4}{2}) - \frac{y^2}{4} |dy+ \int_{0}^{4} |(\frac{y+4}{2}) - \frac{y^2}{4} |dy"

"A = | \frac{1}{2}(\frac{y^2}{2}+4y)-\frac{1}{4}\frac{y^3}{3} |_{-2}^{0} + | \frac{1}{2}(\frac{y^2}{2}+4y)-\frac{1}{4}\frac{y^3}{3} |_{0}^{4}"

Putting limits,

"A = | \frac{1}{2}(\frac{4}{2}-8)-\frac{1}{4}\frac{(-8)}{3} | + | \frac{1}{2}(\frac{16}{2}+16)-\frac{1}{4}\frac{64}{3} | = |\frac{14}{6}|+|\frac{20}{3}| = \frac{54}{6}" square units.